OpenStudy (fanduekisses):

What does the 45' mean in this question???

OpenStudy (fanduekisses):

I have to convert from degrees to radians, I know how to do that but what does the 45' mean?

OpenStudy (smilie2015):

. The 45' means 45/60 of a degree, since each degree contains sixty minutes. Simplification and long division gives me 45/60 = 3/4 = 0.75. So the 45' is 0.75°.

OpenStudy (fanduekisses):

ohhh ok thanks